If an astronaut throws a tool with a force of 16.0N , what is the magnitude of the acceleration a of the astronaut during the throw? Assume that the total mass of the astronaut after she throws the tool is 80.0kg .

(Express your answer in meters per second squared)
Answer #1
Concepts and reason

The concept of Newton’s second and third law is required to answer the problem

First, use Newton’s second law and determine the expression for the acceleration of the astronaut. Then determine the force acting on the astronaut using the Newton’s third law. Finally, use the value of mass and Force acting on the astronaut to determine the magnitude of the acceleration of the astronaut.

Fundamentals

According to Newton’s second law, the net force acting on an object s equal to the product of its mass and acceleration, and it is expressed as,

F=maF = ma

Here, m is the mass of the object and a is the acceleration of the object.

According to Newton’s third law, for every action there is an equal and opposite reaction.

The force acting on the astronaut is given as,

F=maF = ma

Here, m is the mass of the astronaut and a is the acceleration of the astronaut.

Rearrange the above equation for a.

a=Fma = \frac{F}{m}

The acceleration of the astronaut is given as,

a=Fma = \frac{F}{m}

Here, F is the force acting on the astronaut.

According to Newton’s third law, the magnitude of the force acting on the astronaut is equal to the force with which the astronaut threw the tool.

Substitute 16.0 N for F and 80.0 kg for a in the above equation and determine the magnitude of acceleration of the astronaut.

a=16.0N80.0kg=0.200m/s2\begin{array}{c}\\a = \frac{{16.0\,{\rm{N}}}}{{80.0{\rm{ kg}}}}\\\\ = 0.200{\rm{ m/}}{{\rm{s}}^2}\\\end{array}

Ans:

The magnitude of the acceleration of the astronaut is 0.200m/s20.200{\rm{ m/}}{{\rm{s}}^2} .
Answer #2

suck my fat fucking nuts buddy